[Brahms-l] Shift report 20040214 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sat Feb 14 2004 - 23:53:11 EST
------------------------------------------------------------
Date: Sat Feb 14 23:53:11 2004, Shift: 16:00-24:00
Supervisor: Eun-Joo Kim, with: 
Log book#57 pages 1 - 1

Shift Summary:
------------------------------------------------------------
1. Beam for 17:39 - 22.18, about 4.5 hrs.

2. Data runs

    Finished FS 4-deg 1/1B 1.99M tr2 at run 10347    
             MRS 45-deg 6B 1.5M  tr3 at run 10358   
 
                  MRS 45-deg 6B     FS 4-deg 1/2B  
   10350           223K tr3          200K tr2
   10351           228K tr3          200K tr2
   10352           186K tr3          182K tr2 
   10358            77K tr3           53k tr2

  Pedestal 10348, TPC 10349

3. No new detector problem.
   DAQ had an error once, and IO/RunServer restarted. 
------------------------
Plan for next shift

We need 40k more tr2 at FS 1/2B.
However I asked MCR to change D1-D5 pol. from B to A.
(no beam! at 11:45pm) 

1. MRS 45-deg 1050A 1M
2.  FS  4-deg 1/1A  3.2M tr2
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 14 23:54:03 2004

This archive was generated by hypermail 2.1.8 : Sat Feb 14 2004 - 23:54:24 EST